# Universal Transient Stability Analysis: A Large Language Model-Enabled Dynamics Prediction Framework

### v1

@@ Abstract
Existing dynamics prediction frameworks for transient stability analysis (TSA) fail to achieve multi-scenario ”universality”—the inherent ability of a single, pre-trained architecture to generalize across diverse operating conditions, unseen faults, and heterogeneous systems. To address this, this paper proposes TSA-LLM, a large language model (LLM)-based universal framework that models multi-variate transient dynamics prediction as a univariate generative task with three key innovations: First, a novel data processing pipeline featuring channel independence decomposition to resolve dimensional heterogeneity, sample-wise normalization to eliminate separate stable/unstable pipelines, and temporal patching for efficient long-sequence modeling; Second, a parameter-efficient freeze-and-finetune strategy that augments the LLM’s architecture with dedicated input embedding and output projection layers while freezing core transformer blocks to preserve generic feature extraction capabilities; Third, a two-stage fine-tuning scheme that combines teacher forcing, which feeds the model ground-truth data during initial training, with scheduled sampling, which gradually shifts to leveraging model-generated predictions, to mitigate cumulative errors in long-horizon iterative prediction. Comprehensive testing demonstrates the framework’s universality, as TSA-LLM trained solely on the New England 39-bus system achieves zero-shot generalization to mixed stability conditions and unseen faults, and matches expert performance on the larger Iceland 189-bus system with only 5% fine-tuning data. This multi-scenario versatility validates a universal framework that eliminates scenario-specific retraining and achieves scalability via large-scale parameters and cross-scenario training data.
@@
Deep learning, data-driven methods, transient stability analysis, dynamic trajectory prediction, pre-trained large language model, iterative prediction.
@@
[label1]organization=Collage of Control Science and Engineering, Zhejiang University,postcode=310027, city=Hangzhou, country=China
@@ 1 Introduction
The increasing penetration of renewable energy sources reduces system inertia and elevates operational uncertainty, posing significant challenges to grid stability [shen2025physics]. Consequently, modern power systems are increasingly susceptible to contingencies, demanding more robust analytical frameworks. A key approach is transient stability analysis (TSA), which evaluates the dynamic behavior of synchronous generators (SGs) following disturbances [ye2024use]. TSA is critical for preventive control and emergency response, enabling grid operators to enhance resilience against cascading failures and ensure reliable operation under evolving operating conditions (OCs). Conventionally, the dynamic behavior of power systems is modeled using nonlinear differential-algebraic equations (DAEs), which are numerically solved via time-domain simulation (TDS) methods. While TDS provides high-fidelity temporal resolution, its computational intensity and scalability challenges limit its online practicality [zhao2022structure]. Alternatively, direct methods assess transient stability by analyzing energy-based criteria without explicitly solving DAEs [qiu2022adaptive]. These methods evaluate system stability by comparing post-fault energy to a predefined critical energy threshold. However, their reliance on simplified system models and inherently conservative critical energy estimates restricts their applicability in real-world scenarios[azman2020unified].
@@
The widespread deployment of phasor measurement units (PMUs) has spurred the development of data-driven TSA. These methods learn the mapping between system measurements and stability-related features, enabling rapid online applications such as real-time stability classification [zhou1994application, gomez2010support, Sun2007An, chen2022interpretable, ren2021interpretable, azman2020unified, zhu2023integrated] and stability margin estimation [su2023online, zhu2019hierarchical, zhu2021networked]. Early research in this domain was dominated by classical machine learning. For instance, artificial neural networks were applied to perform rapid transient stability classification [zhou1994application], while support vector machines demonstrated robust post-fault stability prediction using synchrophasor data [gomez2010support]. Concurrently, decision trees were used to extract critical stability indicators through offline feature ranking, facilitating periodic updates to security rules [Sun2007An]. More recent research has shifted toward deep learning (DL) frameworks to improve both prediction accuracy and model interpretability. To enhance explainability, a time-adaptive attention-based gated recurrent unit (GRU) was developed to visualize feature importance [chen2022interpretable], while other work integrated decision tree path length constraints into the DL loss function [ren2021interpretable]. For performance improvement, hybrid architectures combining convolutional neural networks (CNNs) and long short-term memory (LSTMs) were proposed for stability classification and oscillation prediction [azman2020unified]. Beyond binary classification, significant effort has focused on directly predicting stability margins, often derived from the critical clearing time (CCT). Notable methods include an ensemble DL framework with dynamic error correction for CCT estimation [su2023online] and a two-level CNN regression model that predicts margins from 2-D images of system trajectories [zhu2019hierarchical].
@@
Although existing data-driven models provide rapid and accurate stability classification, this binary output offers limited actionable insight for system operators [zhao2022structure, ye2024use]. Consequently, TSA research has shifted toward predicting detailed dynamic trajectories, which are critical for informing corrective actions such as load shedding or out-of-step (OOS) SG tripping [ye2024use]. One prominent approach involves physics-informed neural networks (PINNs), which incorporate physical constraints as regularization terms to predict system dynamics [misyris2020physics]. However, the applicability of early PINN models was often limited to stable OCs. Subsequent work on physics-following neural networks (PFNNs) has sought to improve generalization across both stable and unstable cases by introducing time-varying algebraic parameters and supervised initialization into the PINN framework [shen2025physics]. Other notable strategies include using LSTMs with separate processing pipeline for stable and unstable samples [ye2024use] and leveraging Fourier neural operators (FNOs) to predict transients in the frequency domain, thereby simplifying the learning task [cui2023frequency].
@@
While the methods in [shen2025physics, ye2024use, cui2023frequency] improve prediction accuracy in mixed stable/unstable OCs, they exhibit poor generalization to unseen faults. To address this, the stochastic variational deep kernel regressor (SVDKR)-based framework in [tan2025bayesian] integrates an OOS detector for classifying SGs as stable or unstable. Dedicated SVDKR predictors are trained per stability class, followed by per-SG error correctors to refine unit-level predictions. Although this improves robustness, the parameter complexity and computational overhead hinder its scalability to large systems with numerous SGs. The deep neural representation (DNR) framework was proposed to mitigate these issues [zhao2022structure], using a unified graph attention and GRU architecture to predict dynamics in a single model. This design enhances performance under unseen faults but suffers from significant cumulative error amplification over long prediction horizons. More importantly, both the SVDKR [tan2025bayesian] and DNR [zhao2022structure] frameworks are designed for a single system. They lack the ability to generalize across heterogeneous power systems with different topologies and SG configurations, necessitating laborious retraining for each new system.
@@
As demonstrated in Table 1, existing frameworks suffer from pronounced performance degradation across diverse OCs, faults and system configurations. This shortcoming motivates the pursuit of universality in TSA models, defined as the intrinsic capability of a single pre-trained architecture to generalize robustly across mixed stable/unstable OCs, unseen fault events, and heterogeneous power systems. Universality enables a paradigm shift in TSA, analogous to the versatility of large language models (LLMs) in natural language processing, where a unified model adeptly addresses diverse dynamic trajectory prediction tasks [achiam2023gpt]. By eliminating the need for scenario-specific retraining, a universal TSA framework provides foundational capabilities, most prominently exemplified by cross-system generalization. This feature enables efficient knowledge transfer from source-domain systems to large-scale target grids via few-shot adaptation, thereby minimizing computational simulation and labor costs associated with custom model development for diverse power networks [tu2024powerpm]. Furthermore, universality inherently supports scalability, with performance systematically improving as model parameters and multi-scenario training datasets expand, which ensures adaptability to evolving grid complexities [liu2024timer].
@@
However, achieving this universality requires overcoming three fundamental limitations: 1) Data processing rigidity, where current methods are constrained to fixed input dimensionalities based on specific state variables, limiting scalability and requiring complete retraining for system changes or monitored SGs modifications [zhao2022structure], while frequently necessitating separate processing pipelines for stable/unstable dynamics that introduce computational overhead and misclassification risks [tan2025bayesian, ye2024use]; 2) Architectural constraints, as prevalent recurrent neural network architectures (LSTMs [ye2024use] and GRUs [zhao2022structure]) suffer from vanishing gradient problems that impede long-range dependency modeling and inherently sequential processing that precludes parallel computation, creating critical bottlenecks for model scalability [liu2024timer]; 3) Training and error propagation, where the temporal asymmetry between short observation windows and extended prediction horizons in online TSA, combined with existing training schemes that lack explicit error suppression mechanisms, causes prediction errors to accumulate at each time step during iterative prediction, severely degrading long-term reliability [shen2025physics, zhao2022structure].
@@
To address these challenges, this study proposes a pre-trained LLM-based universal dynamics prediction framework, TSA-LLM, that formulates multi-variate transient dynamics prediction as a univariate generative sequence prediction task. The main contributions are:
@@
    To the best of the authors’ knowledge, this is the first work to demonstrate universality in TSA through an LLM-based approach, achieving generalization across diverse OCs, unseen faults, and heterogeneous systems by systematically addressing fundamental limitations in data processing, model architecture, and fine-tuning schemes.
@@
    A novel data processing pipeline is designed and operates sequentially: channel-independence decouples multivariate dynamic sequences into univariate streams, resolving dimensional heterogeneity across different systems; sample-wise normalization standardizes sequence samples via internal statistics, eliminating separate pipelines for stable/unstable OCs; temporal patching tokenizes sequences into semantic patches, enabling efficient long-sequence modeling while capturing high-level temporal patterns.
@@
    A Generative Pre-trained Transformer (GPT) architecture is devised that incorporates dedicated input embedding and output projection alongside causal attention mechanisms for temporal dependency modeling. The freeze-and-finetune strategy preserves pre-trained sequence feature extraction capability by freezing core transformer blocks (T-blocks) while selectively fine-tuning TSA-specific layers, achieving balance between knowledge retention and task adaptation.
@@
    A two-stage fine-tuning scheme is proposed to mitigate error propagation during iterative and generative dynamics prediction. The initial teacher forcing (TeaF) stage ensures stable convergence by exclusively feeding ground-truth sequences as input. Subsequently, the scheduled sampling (SchS) stage enhances robustness by increasingly replacing ground-truth inputs with the model’s own predictions, thereby compelling the model to correct for self-generated errors.
@@
Method Universality Scenarios
Single
Model
Stable/Unstable
OCs
Unseen
Faults
Heterogeneous
System
PINN [misyris2020physics] $\times$ $\times$ $\times$ $\checkmark$
LSTM [ye2024use] $\checkmark$ $\times$ $\times$ $\times$
PFNN [shen2025physics] $\checkmark$ $\times$ $\times$ $\checkmark$
FNO [cui2023frequency] $\checkmark$ $\times$ $\times$ $\checkmark$
SVDKR [tan2025bayesian] $\checkmark$ $\checkmark$ $\times$ $\times$
DNR [zhao2022structure] $\checkmark$ $\checkmark$ $\times$ $\checkmark$
@@ 2.1 Power System Model
Considering a power system containing $n_{b}$ buses, $n_{g}$ SGs, and $n_{l}$ loads, the DAEs describing the the system’s transient dynamics, can be formulated as follows [shen2025physics]:

### v2

# Universal Transient Stability Analysis: A Large Language Model-Enabled Power System Dynamics Prediction Framework

@@ Abstract
Existing dynamics pblackiction frameworks for transient stability analysis (TSA) fail to achieve multi-scenario ”universality”—the inherent ability of a single, pre-trained architecture to generalize across diverse operating conditions, unseen faults, and heterogeneous systems. To address this, this paper proposes TSA-LLM, a large language model (LLM)-based universal framework that models multivariate transient dynamics pblackiction as a univariate generative task with three key innovations: First, a novel data processing pipeline featuring channel independence decomposition to resolve dimensional heterogeneity, sample-wise normalization to eliminate separate stable/unstable pipelines, and temporal patching for efficient long-sequence modeling; Second, a parameter-efficient freeze-and-finetune strategy that augments the LLM’s architecture with dedicated input embedding and output projection layers while freezing core transformer blocks to preserve generic feature extraction capabilities; Third, a two-stage fine-tuning scheme that combines teacher forcing, which feeds the model ground-truth data during initial training, with scheduled sampling, which gradually shifts to leveraging model-generated pblackictions, to mitigate cumulative errors in long-horizon iterative pblackiction. Comprehensive testing demonstrates the framework’s universality, as TSA-LLM trained solely on the New England 39-bus system achieves zero-shot generalization to mixed stability conditions and unseen faults, and matches expert performance on the Iceland 189-bus system with only 5% fine-tuning data. Additional cross-system experiments on the IEEE 68-bus and IEEE 118-bus systems, together with stability metrics and PEBS comparison, further confirm TSA-LLM’s strong zero-shot transferability and data-efficient adaptation. This multi-scenario versatility validates a universal framework that eliminates scenario-specific retraining and achieves scalability via large-scale parameters and cross-scenario training data.
@@
Deep learning, data-driven methods, transient stability analysis, dynamic trajectory pblackiction, pre-trained large language model, iterative pblackiction.
@@
[label1]organization=College of Control Science and Engineering, Zhejiang University,postcode=310027, city=Hangzhou, country=China
@@ 1 Introduction
The increasing penetration of renewable energy sources blackuces system inertia and elevates operational uncertainty, posing significant challenges to grid stability [29, 25]. Consequently, modern power systems are increasingly susceptible to contingencies, demanding more robust analytical frameworks. Power system stability is the ability of an electric power system, for a given initial operating condition, to regain a state of operating equilibrium after being subjected to a physical disturbance, with most system variables bounded so that practically the entire system remains intact [12]. A key aspect of this broad stability notion is transient stability analysis (TSA), which assesses whether the system can maintain synchronism and reach an acceptable post-disturbance equilibrium following a large disturbance (e.g., short-circuit faults, line/generator tripping), as manifested in the post-disturbance transient dynamics [30, 40, 12]. Accordingly, the core objective of this work is to support TSA by rapidly pblackicting post-fault trajectories of key state variables (e.g., rotor angle and speed) from limited PMU observations, thereby enabling timely stability assessment and emergency decision-making.
TSA is critical for preventive control and emergency response, enabling grid operators to enhance resilience against cascading failures and ensure reliable operation under evolving operating conditions (OCs) [44]. Conventionally, the dynamic behavior of power systems is typically described by nonlinear differential-algebraic equations (DAEs), which are numerically solved via time-domain simulation (TDS) methods [42]. While TDS provides high-fidelity temporal resolution, its computational intensity and scalability challenges limit its online practicality [43]. Alternatively, direct methods evaluate transient stability without performing long-horizon time-domain simulation. Following the taxonomy summarized in [22], mainstream direct methods can be categorized into three groups. (i) EAC-based equivalencing methods (e.g., the Extended Equal Area Criterion (EEAC) [39]) extend the equal-area criterion from SMIB to multi-machine systems by coherent-grouping and two-machine equivalencing [15]. In practice, they often become hybrid integral-direct methods due to the need for trajectory-dependent grouping or auxiliary integration. (ii) Global energy-function (topological) methods [5] construct a global energy function $V(x)$ and characterize the stability region via the closest controlling unstable equilibrium point (UEP) [3]. A representative practical variant is the Potential Energy Boundary Surface (PEBS) method, which approximates the stability boundary by locating the first local maximum of the potential energy along the fault-on trajectory, thereby avoiding explicit UEP computation [7]. While these methods provide clear geometric interpretation and stability margins, they rely on strong model assumptions (existence of a global energy function) and typically still require obtaining the fault-clearing state. (iii) Local Lyapunov-function (LLF) methods [22, 50, 38] construct a Lyapunov function locally around the post-fault stable equilibrium and certify stability via an invariant sublevel set, which can, in principle, avoid numerical integration and UEP computation. However, traditional LLF conditions and set/derivative bounds may introduce conservatism. Despite their theoretical elegance, direct methods face challenges in model generality, numerical tractability, and conservatism in large-scale online settings [22, 2].
@@
The widespread deployment of phasor measurement units (PMUs) has spurblack the development of data-driven TSA. These methods learn the mapping between system measurements and stability-related features, enabling rapid online applications such as real-time stability classification [45, 11, 33, 6, 24, 2, 49] and stability margin estimation [32, 47, 48]. Early research in this domain was dominated by classical machine learning. For instance, artificial neural networks were applied to perform rapid transient stability classification [45], while support vector machines demonstrated robust post-fault stability pblackiction using synchrophasor data [51, 11]. Concurrently, decision trees were used to extract critical stability indicators through offline feature ranking, facilitating periodic updates to security rules [17, 33]. More recent research has shifted toward deep learning (DL) frameworks to improve both pblackiction accuracy and model interpretability. To enhance explainability, a time-adaptive attention-based gated recurrent unit (GRU) was developed to visualize feature importance [6], while other work integrated decision tree path length constraints into the DL loss function [24]. For performance improvement, hybrid architectures combining convolutional neural networks (CNNs) and long short-term memory (LSTMs) were proposed for stability classification and oscillation pblackiction [2, 30]. Beyond binary classification, significant effort has focused on directly pblackicting stability margins, often derived from the critical clearing time (CCT). Notable methods include an ensemble DL framework with dynamic error correction for CCT estimation [32] and a two-level CNN regression model that pblackicts margins from 2-D images of system trajectories [47].
@@
Although existing data-driven models provide rapid and accurate stability classification, this binary output offers limited actionable insight for system operators [43, 40]. Consequently, TSA research has shifted toward pblackicting detailed dynamic trajectories, which are critical for informing corrective actions such as load shedding or out-of-step (OOS) SG tripping [40]. One prominent approach involves physics-informed neural networks (PINNs), which incorporate physical constraints as regularization terms to pblackict system dynamics [21]. However, the applicability of early PINN models was often limited to stable OCs. Subsequent work on physics-following neural networks (PFNNs) has sought to improve generalization across both stable and unstable cases by introducing time-varying algebraic parameters and supervised initialization into the PINN framework [29]. Other notable strategies include using LSTMs with separate processing pipeline for stable and unstable samples [40] and leveraging Fourier neural operators (FNOs) to pblackict transients in the frequency domain, thereby simplifying the learning task [8].
@@
While the methods in [29, 40, 8] improve pblackiction accuracy in mixed stable/unstable OCs, they exhibit poor generalization to unseen faults. To address this, the stochastic variational deep kernel regressor (SVDKR)-based framework in [34] integrates an OOS detector for classifying SGs as stable or unstable. Dedicated SVDKR pblackictors are trained per stability class, followed by per-SG error correctors to refine unit-level pblackictions. Although this improves robustness, the parameter complexity and computational overhead hinder its scalability to large systems with numerous SGs. The deep neural representation (DNR) framework was proposed to mitigate these issues [43], using a unified graph attention and GRU architecture to pblackict dynamics in a single model. This design enhances performance under unseen faults but suffers from significant cumulative error amplification over long pblackiction horizons. More importantly, both the SVDKR [34] and DNR [43] frameworks are designed for a single system. They lack the ability to generalize across heterogeneous power systems with different topologies and SG configurations, necessitating laborious retraining for each new system. This fundamental limitation—the absence of cross-system generalization capability—represents a critical knowledge gap in current TSA research, preventing the development of universal frameworks that could enable efficient knowledge transfer across diverse power networks.
@@
As demonstrated in Table 2, existing frameworks suffer from pronounced performance degradation across diverse OCs, faults and system configurations. This systematic failure across multiple dimensions reveals a fundamental knowledge gap: no existing TSA framework has achieved true universality. This shortcoming motivates the pursuit of universality in TSA models, defined as a novel paradigm representing the intrinsic capability of a single pre-trained architecture to generalize robustly across mixed stable/unstable OCs, unseen fault events, and heterogeneous power systems. Universality enables a paradigm shift in TSA, analogous to the versatility of large language models (LLMs) in natural language processing, where a unified model adeptly addresses diverse dynamic trajectory pblackiction tasks [1]. By eliminating the need for scenario-specific retraining, a universal TSA framework provides foundational capabilities, most prominently exemplified by cross-system generalization. This feature enables efficient knowledge transfer from source-domain systems to large-scale target grids via few-shot adaptation, thereby minimizing computational simulation and labor costs associated with custom model development for diverse power networks [35]. Furthermore, universality inherently supports scalability, with performance systematically improving as model parameters and multi-scenario training datasets expand, which ensures adaptability to evolving grid complexities [19].
To avoid ambiguity, this paper adopts an operational definition of TSA “universality” along three explicit generalization axes: (i) operating-condition universality, i.e., robustness to mixed stable/unstable post-fault trajectories under diverse pre-fault operating points; (ii) fault universality, i.e., generalization to unseen fault locations/clearing conditions beyond those observed in training; and (iii) cross-system universality, i.e., reusing the same pre-trained architecture and model interface across heterogeneous networks with different SG configurations, with at most lightweight few-shot adaptation instead of full retraining.
@@
However, achieving this universality requires overcoming three fundamental limitations: 1) Data processing rigidity, where current methods are constrained to fixed input dimensionalities based on specific state variables, limiting scalability and requiring complete retraining for system changes or monitoblack SGs modifications [43], while frequently necessitating separate processing pipelines for stable/unstable dynamics that introduce computational overhead and misclassification risks [34, 40]; More importantly for cross-system generalization, the number of monitoblack SGs typically differs across networks, which directly changes the dimension of system-level stacked trajectories. This creates an inherent input/output dimension mismatch for pblackictors with fixed multivariate interfaces, and is a key reason why many existing models must be blackefined or retrained when transferring to a new system. 2) Architectural constraints, as prevalent recurrent neural network architectures (LSTMs [40] and GRUs [43]) suffer from vanishing gradient problems that impede long-range dependency modeling and inherently sequential processing that precludes parallel computation, creating critical bottlenecks for model scalability [9, 19]; 3) Training and error propagation, where the temporal asymmetry between short observation windows and extended pblackiction horizons in online TSA, combined with existing training schemes that lack explicit error suppression mechanisms, causes pblackiction errors to accumulate at each time step during iterative pblackiction, severely degrading long-term reliability [29, 13].
@@
To address these challenges, this study proposes a pre-trained LLM-based universal dynamics pblackiction framework, TSA-LLM, that fundamentally reformulates multivariate transient dynamics pblackiction as a univariate generative sequence pblackiction task by leveraging pre-trained LLM parameters whose modality-agnostic sequence feature extraction and temporal dependency modeling capabilities transfer effectively to power-system dynamics (see A for empirical validation). The main contributions are:
@@
    To the best of the authors’ knowledge, this is the first work to demonstrate universality in TSA through an LLM-based approach, achieving generalization across diverse OCs, unseen faults, and heterogeneous systems by systematically addressing fundamental limitations in data processing, model architecture, and fine-tuning schemes. This work establishes a new research direction by bridging generative sequence modeling and power system dynamics pblackiction. The challenge-to-design mapping is summarized in Table 1.
@@
    A novel data processing pipeline is designed and operates sequentially: channel-independence decouples multivariate dynamic sequences into univariate streams, resolving dimensional heterogeneity across different systems and enabling system-agnostic processing; sample-wise normalization standardizes sequence samples via internal statistics, eliminating separate pipelines for stable/unstable OCs; temporal patching tokenizes sequences into semantic patches, enabling efficient long-sequence modeling while capturing high-level temporal patterns. This pipeline design transforms the problem from multivariate regression to univariate sequence generation, enabling direct compatibility with pre-trained LLM architectures.
@@
    A Generative Pre-trained Transformer (GPT) architecture is devised that incorporates dedicated input embedding and output projection alongside causal attention mechanisms for temporal dependency modeling. This represents the first application of pre-trained transformer architectures to power system dynamics pblackiction, leveraging transfer learning from large-scale sequence modeling tasks. The freeze-and-finetune strategy preserves pre-trained sequence feature extraction capability by freezing core transformer blocks (T-blocks) while selectively fine-tuning TSA-specific layers, achieving balance between knowledge retention and task adaptation.
@@
    A two-stage fine-tuning scheme is proposed to mitigate error propagation during iterative and generative dynamics pblackiction. This training strategy is specifically designed for autoregressive sequence generation in TSA, addressing the unique challenge of error accumulation in long-horizon pblackiction tasks. The initial teacher forcing (TeaF) stage ensures stable convergence by exclusively feeding ground-truth sequences as input. Subsequently, the scheduled sampling (SchS) stage enhances robustness by increasingly replacing ground-truth inputs with the model’s own pblackictions, thereby compelling the model to correct for self-generated errors.
Table 1: Challenge-to-design mapping
Challenge Design solution
Data processing rigidity Channel-independence, sample-wise normalization, temporal patching (Sec. 3.2)
Architectural constraints GPT with causal attention, freeze-and-finetune strategy (Secs. 3.2–3.3)
[…]
@@
Method Universality Scenarios Single Model
Stable/Unstable OCs Unseen Faults Heterogeneous System
PINN [21] $\times$ $\times$ $\times$ $\checkmark$
LSTM [40] $\checkmark$ $\times$ $\times$ $\times$
PFNN [29] $\checkmark$ $\times$ $\times$ $\checkmark$
FNO [8] $\checkmark$ $\times$ $\times$ $\checkmark$
SVDKR [34] $\checkmark$ $\checkmark$ $\times$ $\times$
DNR [43] $\checkmark$ $\checkmark$ $\times$ $\checkmark$
@@ 2.1 Power System Model
Considering a power system containing $n_{b}$ buses, $n_{g}$ SGs, and $n_{l}$ loads, the DAEs describing the system’s transient dynamics can be formulated as follows [29, 37]: